# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ 3 Chowla’s cosine problem
In this section, we give a short self–contained proof of Theorem 1.1, assuming Theorem 1.2. We begin the section by recalling some standard notation. Let $G$ be a finite group, $A\subset G$ be a symmetric subset (i.e. a set satisfying $A=A^{-1}$), and let $\Gamma=\mbox{Cay}(G,A)$. Recall that $\mbox{Cay}(G,A)$ is the Cayley graph on $G$ generated by $A$, that is, the graph on vertex set $G$ in which $x,y\in G$ are joined by an edge if $xy^{-1}\in A$. In case $G$ is abelian, we use ++ to denote the group operation. In this case, it is well known that the eigenvalues of $G$ are the values of the discrete Fourier transform $\widehat{\mathds{1}_{A}}$. In the special case $G=\mathbb{Z}/n\mathbb{Z}$, this gives that the eigenvalues of the Cayley graph are
@@
There exists an absolute constant $\gamma>0$ such that for any set $A$ of positive integers, there exists $x\in[0,2\pi]$ such that
$\sum_{a\in A}\cos(ax)\leq-\Omega(|A|^{\gamma}).$
@@
Without loss of generality, we may assume $|A|$ is sufficiently large. Let $n>4|A|$ be a prime larger than all elements of $A$, and let $\Gamma=\mbox{Cay}(\mathbb{Z}/n\mathbb{Z},A\cup-A)$ be the Cayley graph with the generating set $A\cup-A$ (where $A=\{-a:a\in A\}$). Then $\Gamma$ is an $n$-vertex $d$-regular graph with $d=2|A|$. Each $\xi\in\mathbb{Z}/n\mathbb{Z}$ corresponds to an eigenvalue of $\Gamma$ given by
$\lambda_{\xi}=\sum_{a\in A\cup-A}e^{2\pi ia\xi/n}=2\sum_{a\in A}\cos\left(\frac{2\pi a\xi}{n}\right)$
for some $\xi\in\mathbb{Z}/n\mathbb{Z}$. Hence, if we denote by $\lambda_{n}$ the smallest eigenvalue of $\Gamma$, there exists $x=\frac{2\pi\xi}{n}$ such that
@@
Thus, our aim is to show that $|\lambda_{n}|\geq d^{\gamma}$, for an absolute constant $\gamma>0$, since this implies that $\sum_{a\in A}\cos(ax)\leq-d^{\gamma}\leq-\Omega(|A|^{\gamma})$, as needed. Assume, for the sake of contradiction, that $|\lambda_{n}|<d^{\gamma}$. We use Theorem 1.2 to find a clique $S$ of size $|S|=d^{1-c\gamma}$ for some absolute constant $c$, assuming $d$ is sufficiently large with respect to $\gamma$. We now argue that $\Gamma$ cannot contain such large cliques.
@@
As $S$ is a clique in $\Gamma$, we have $S-S\subset A\cup-A\cup\{0\}$. By a simple averaging argument, there exists $t\in A\cup-A$ such that $s-s^{\prime}=t$ for at least $\frac{|S|(|S|-1)}{2|A|}=\frac{|S|(|S|-1)}{d}$ pairs $(s,s^{\prime})\in S\times S$. Hence, for at least |S|(|S|1)/d|S|(|S|-1)/d values of $s\in S$ we have $s+t\in S$, and therefore $|(t+S)\cap S|\geq|S|(|S|-1)/d$. ∎
For a positive integer $k$, let $H_{k}$ is the graph that is formed by a clique of size $2k$, and an additional vertex connected to half of the vertices of the clique.
@@
Let $X\cup Y$ be the partition of $C$, where $X$ is the neighbourhood of $x_{0}$. Let $B$ be the adjacency matrix of $H_{k}$ and let $v\in\mathbb{R}^{V(H_{k})}$ be the vector defined as $v(x_{0})=1/\sqrt{2}$, $v(y)=-\frac{1}{2\sqrt{k}}$ if $y\in X$, and $v(y)=\frac{1}{2\sqrt{k}}$ if $y\in Y$. Then $v||_{2}=1$ and
$v^{T}Bv=2\sum_{ab\in E(H_{k})}v(a)v(b)=2k\cdot\frac{1}{\sqrt{2}}\cdot\big{(}-\frac{1}{2\sqrt{k}}\big{)}+4\cdot\binom{k}{2}\cdot\frac{1}{4k}-2\cdot k^{2}\cdot\frac{1}{4k}=-\sqrt{\frac{{k}}{2}}-\frac{1}{2}<-\sqrt{\frac{k}{2}},$
where the first term is the contribution of the edges $x_{0}y$, $y\in X$, the second term is the contribution of edges $yy^{\prime}$ with $y,y^{\prime}\in X$ or $y,y^{\prime}\in Y$, and the third term is the contribution of edges $yy^{\prime}$ with $y\in X$ and $y^{\prime}\in Y$. ∎
By Claim 3.3, the smallest eigenvalue $\mu$ of $H_{k}$ satisfies $\mu<-\frac{1}{2}\sqrt{k}$. Therefore, by the Cauchy interlacing theorem, $\Gamma$ does not contain $H_{k}$ as an induced subgraph for $k=4d^{2\gamma}$. As $|S|>2k$, each vertex of $\Gamma$ sends either at most $k$ edges to $S$, or at least |S|k|S|-k edges. We prove that every vertex in $\Gamma$ must send at least |S|k|S|-k edges to $S$. This easily leads to a contradiction for $n$ sufficiently large: this implies that there are at least $(n-|S|)(|S|-k)\geq\frac{n}{2}\cdot\frac{|S|}{2}>d|S|$ edges with an endpoint in $S$, contradicting that $\Gamma$ is $d$-regular.
@@
Every $v\in V(\Gamma)$ sends at least |S|k|S|-k edges to $S$.
@@
We prove by induction on $\ell$ that every vertex of $\ell t+S$ sends at least |S|k|S|-k edges to $S$. As every vertex $v\in V(\Gamma)$ is contained in some $\ell t+S$, this finishes the proof. The base case $\ell=0$ is trivial, so let $\ell\geq 1$. By our induction hypothesis and translation invariance, every vertex $v\in\ell t+S$ sends at least |S|k|S|-k edges to $t+S$. But then $v$ sends at least
$|S\cap(t+S)|-k=\Omega(|S|^{2}/d)-k=\Omega(d^{1-2c\gamma})-4d^{2\gamma}>4d^{2\gamma}=k$
edges to $S\cap(t+S)$, and in particular, more than $k$ edges to $S$. Here the last inequality holds if we take $\gamma=\frac{1}{3+2c}$ and $|A|$ sufficiently large. Therefore, as $\Gamma$ contains no induced copy of $H_{k}$, $v$ must send at least |S|k|S|-k edges to $S$, and we are done. ∎
@@ 4 Preliminaries
We recall some basic facts and standard notation from linear algebra and graph theory. The edge density of a graph $G$ is $m/\binom{n}{2}$, where $m=e(G)$ is the number of edges. Given a subset $U$ of the vertices, $G[U]$ denotes the subgraph of $G$ induced on vertex set $U$. Also, if $V\subset V(G)$ is disjoint form $U$, then $G[U,V]$ is the bipartite subgraph of $V(G)$ induced between $U$ and $V$. The complement of $G$ is denoted by $\overline{G}$. The maximum degree of $G$ is denoted by $\Delta(G)$, and the average degree by $d(G)$. If $G$ has $n$ vertices, we will often identify the set of vertices of $G$ with $[n]$.
@@
Given an $n\times n$ real symmetric matrix $M$, we denote by $\lambda_{1}(M)\geq\dots\geq\lambda_{n}(M)$ the eigenvalues of $M$ with multiplicity. As we will see, the surplus of $G$ is controlled by the negative eigenvalues of the graph. If $G$ is an $n$-vertex graph whose adjacency matrix is $A$, then we denote by $\lambda_{i}=\lambda_{i}(A)$ the eigenvalues of $A$, sometimes also calling them the eigenvalues of $G$. We also denote by $v_{1},\dots,v_{n}$ a corresponding orthonormal basis of eigenvectors. By the Perron–Frobenius theorem, we may take $v_{1}$ to be a vector with non-negative entries, which we call the principal eigenvector of $A$. Furthermore, the corresponding eigenvalue satisfies $\lambda_{1}\geq d(G)$. See the survey [24] as a general reference on the principal eigenvector.
@@
We also use the Hadamard product for vectors: for $u,v\in\mathbb{R}^{n}$, their Hadamard product vector $u\circ v\in\mathbb{R}^{n}$ is defined by $(u\circ v)(i):=u(i)v(i)$ for all $i\in\{1,\dots,n\}$.
@@ 5 Spectral lower bounds for the surplus
In this section, we present bounds on the MaxCut of a graph in terms of its spectrum. These inequalities will be crucial in transferring our results for the smallest eigenvalue to the MaxCut setting.
@@
For an $n$-vertex graph $G$ with the smallest eigenvalue $\lambda_{n}$, we have $\operatorname{surp}(G)\leq|\lambda_{n}|n/4$.
@@
But, we have $x^{T}Ax\leq|\lambda_{n}|\|x\|_{2}^{2}$ for every vector $x\in\mathbb{R}^{n}$, and so $\operatorname{surp}(G)\leq\frac{1}{4}|\lambda_{n}|\sqrt{n}^{2}=|\lambda_{n}|n/4$. ∎
The key ingredient of the above proof is the relation $\operatorname{surp}(G)=\frac{1}{4}\max_{x\in[-1,1]^{n}}-x^{T}Ax$, which can also be written as $\operatorname{surp}(G)=\frac{1}{4}\max_{x\in[-1,1]^{n}}\langle-A,xx^{T}\rangle$, where we observe that $xx^{T}$ is a positive-semidefinite matrix with diagonal entries bounded by $1$. As we will see, it will be very convenient to define the semidefinite relaxation of the surplus as follows. Given an $n$-vertex graph $G$ with adjacency matrix $A$, define
@@
$v||_{\infty}\leq\frac{\sqrt{n}}{|\lambda|}.$
@@
where we used that $\sum_{i=1}^{n}v_{1}(i)^{2}=1$. Hence $|\lambda||v(b)|\leq\sqrt{n}$. ∎
@@
Let $G$ be an $n$-vertex graph, whose complement has edge density $p\leq 1/10$ and maximum degree $\bar{\Delta}=\Delta(\overline{G})$. If $v_{1}$ is the principal eigenvector of $G$, then for each $i\in[n]$ we have
$\frac{1-2\bar{\Delta}/n}{\sqrt{n}}\leq v_{1}(i)\leq\frac{1+2p+2/n}{\sqrt{n}}.$
@@
Let d=d(G)=(1p)(n1)d=d(G)=(1-p)(n-1) be the average degree of $G$, and recall that $\lambda_{1}\geq d$. By Lemma 5.4,
$v_{1}(b)\leq\frac{\sqrt{n}}{\lambda_{1}}\leq\frac{\sqrt{n}}{d}=\frac{\sqrt{n}}{(1-p)(n-1)}\leq\frac{1+2p+2/n}{\sqrt{n}}.$
In the last inequality, we used that $p<1/10$. To prove the lower bound, we first observe that
$1=\sum_{i=1}^{n}v_{1}(i)^{2}\leq\|v_{1}\|_{\infty}\sum_{i=1}^{n}v_{1}(i),$
which implies that $\sum_{i=1}^{n}v_{1}(i)\geq\frac{\lambda_{1}}{\sqrt{n}}$. But then using the identity $Av_{1}=\lambda_{1}v_{1}$,
$\lambda_{1}v_{1}(b)=\sum_{i\sim b}v_{1}(i)\geq\sum_{i=1}^{n}v_{1}(i)-\bar{\Delta}\|v_{1}\|_{\infty}\geq\frac{\lambda_{1}}{\sqrt{n}}-\bar{\Delta}\frac{\sqrt{n}}{\lambda_{1}}=\frac{\lambda_{1}}{\sqrt{n}}\Big{(}1-\bar{\Delta}\frac{n}{\lambda_{1}^{2}}\Big{)}\geq\frac{\lambda_{1}}{\sqrt{n}}\Big{(}1-\frac{2\bar{\Delta}}{n}\Big{)},$
where we used that $\lambda_{1}^{2}\geq d^{2}\geq n^{2}/2$ in the last inequality. Canceling $\lambda_{1}$ gives $v_{1}(b)\geq\frac{1-2\bar{\Delta}/n}{\sqrt{n}}$. ∎
Proof of Lemma 5.3..
@@
(i) Let $X=\sum_{\lambda_{i}<0}v_{i}v_{i}^{T}$. Then $X$ is positive semidefinite, and as $v_{1},\dots,v_{n}$ is an orthonormal basis, we have
@@
(iii) Let $\beta=\frac{1}{100(\Delta^{*}+1)}$, and $X=\beta\sum_{\lambda_{i}<0}\lambda_{i}^{2}v_{i}v_{i}^{T}$. Then $X$ is positive semidefinite. It is enough to prove that the diagonal entries of $X$ are bounded by $1$, as then $\operatorname{surp}^{*}(G)\geq-\langle A,X\rangle=\beta\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}$.
First, we consider the (easier) case $\Delta^{*}=\Delta(G)$. We observe that
@@
is positive semidefinite, so the diagonal entries of $\beta A^{2}$ dominate the diagonal entries of $X$. But $(A^{2})_{i,i}$ is the degree of vertex $i$, so $(A^{2})_{i,i}\leq\Delta\leq 1/\beta$, and the claim follows.
Next, we consider the case $\Delta^{*}=\Delta(\overline{G})$. We may assume that the edge density of $\overline{G}$ is less than $1/10$, otherwise $\Delta^{*}=\Omega(n)$, and the previous case implies $\operatorname{surp}^{*}(G)=\Omega(\frac{1}{n})\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}$. To show that $X_{i,i}\leq 1$, we analyze the matrix $B=A-\lambda_{1}v_{1}v_{1}^{T}$. Since
@@
we have that $\beta B^{2}-X$ is positive semidefinite. Hence, $(\beta B^{2})_{i,i}\geq X_{i,i}$ for every $i\in[n]$. Therefore, it is enough to show that $(B^{2})_{i,i}\leq 1/\beta=100(\Delta^{*}+1)$.
To show this, we first bound the entries of $B$. We denote by $p$ the density of $\overline{G}$, and observe that $p\leq\frac{\Delta^{*}n/2}{\binom{n}{2}}=\frac{\Delta^{*}}{n-1}$. Then, Lemma 5.5 implies that for any $i,j\in[n]$ we have
$1-\frac{5(\Delta^{*}+1)}{n}\leq(n-1)(1-p)\left(\frac{1-2\Delta^{*}/n}{\sqrt{n}}\right)^{2}\leq\lambda_{1}v_{1}(i)v_{1}(j)\leq n\left(\frac{1+2p+2/n}{\sqrt{n}}\right)^{2}\leq 1+\frac{5(\Delta^{*}+1)}{n}.$
Therefore, for every $i,j\in[n]$, if $ij\in E(G)$ and $A_{i,j}=1$, then $|B_{i,j}|\leq\frac{5(\Delta^{*}+1)}{n}$. Otherwise, we have $|B_{i,j}|\leq 1+5(\Delta^{*}+1)/n\leq 6$. From this, we have
$(B^{2})_{i,i}=\sum_{j=1}^{n}(B_{i,j})^{2}\leq 36\Delta^{*}+n\frac{25(\Delta^{*}+1)^{2}}{n^{2}}\leq 100(\Delta^{*}+1).$
@@
Note that the first inequality is the combination of (i) and (iii), while the second one is simply the Cauchy-Schwartz inequality applied to the sequences $(|\lambda_{i}|^{3})_{\lambda_{i}<0}$ and $(|\lambda_{i}|)_{\lambda_{i}<0}$. Taking square roots then proves (ii). ∎
@@
Let $A^{\prime}$ be the adjacency matrix of $G^{\prime}$ and let $X^{\prime}\in\mathbb{R}^{V(G^{\prime})\times V(G^{\prime})}$ be a matrix such that $X^{\prime}$ is positive semidefinite, $X_{i,i}\leq 1$ for every $i\in V(G^{\prime})$, and $\operatorname{surp}^{*}(G^{\prime})=-\langle A^{\prime},X^{\prime}\rangle$. Let $X\in\mathbb{R}^{V(G)\times V(G)}$ be the matrix that agrees with $X^{\prime}$ on every entry $(x,y)\in V(G^{\prime})\times V(G^{\prime})$, and zero everywhere else. Then
$\operatorname{surp}^{*}(G)\geq-\langle A,X\rangle=-\langle A^{\prime},X^{\prime}\rangle=\operatorname{surp}^{*}(G^{\prime}).$

### v2

@@ 3 Chowla’s cosine problem
In this section, we give a short proof of Theorem˜1.1, assuming Theorem 1.2. We begin the section by recalling some standard notation. Let $\Gamma$ be a finite group, $A\subset\Gamma$ be a symmetric subset (i.e. a set satisfying $A=A^{-1}$), and let $G=\operatorname{Cay}(\Gamma,A)$. Recall that $\operatorname{Cay}(\Gamma,A)$ is the Cayley graph on $\Gamma$ generated by $A$, that is, the graph on vertex set $\Gamma$ in which $x,y\in\Gamma$ are joined by an edge if $xy^{-1}\in A$. If $\Gamma$ is abelian, it is well known that the eigenvalues of $G$ are the values of the discrete Fourier transform $\widehat{\mathds{1}_{A}}$. In the special case $\Gamma=\mathbb{Z}/n\mathbb{Z}$, this gives that the eigenvalues of the Cayley graph are
@@
For any finite set $A$ of positive integers, there exists $x\in[0,2\pi]$ such that
$\sum_{a\in A}\cos(ax)\leq-\Omega(|A|^{1/10-o(1)}).$
@@
Let $n>4\max A$ be a prime, and let $G=\operatorname{Cay}(\mathbb{Z}/n\mathbb{Z},A\cup-A)$. Then $G$ is an $n$-vertex $d$-regular graph with $d=2|A|$. Every $\xi\in\mathbb{Z}/n\mathbb{Z}$ corresponds to an eigenvalue of $G$ given by
$\lambda_{\xi}=\sum_{a\in A\cup-A}e^{2\pi ia\xi/n}=2\sum_{a\in A}\cos\left(\frac{2\pi a\xi}{n}\right).$
Hence, if we denote by $\lambda_{n}$ the smallest eigenvalue of $G$, then there exists $x=\frac{2\pi\xi}{n}$ such that
@@
Let $\gamma=1/10-\varepsilon$ for any fixed $\varepsilon>0$. Our aim is to show that $|\lambda_{n}|\geq d^{\gamma}$ for $d$ sufficiently large. Assume to the contrary that $|\lambda_{n}|<d^{\gamma}$, then by Theorem 1.2, $G$ contains a clique $S$ of size $|S|\geq d^{1-4\gamma}=d^{3/5+4\varepsilon}$. We now argue that this is impossible with the help of two auxiliary claims.
@@
As $S$ is a clique in $G$, we have $S-S\subset A\cup-A\cup\{0\}$. By averaging, there exists some $t\in A\cup-A$ such that $s^{\prime}-s=t$ holds for at least $\frac{|S|(|S|-1)}{2|A|}=\frac{|S|(|S|-1)}{d}$ pairs $(s,s^{\prime})\in S\times S$. Hence, for at least |S|(|S|1)/d|S|(|S|-1)/d values of $s\in S$ we have $s+t\in S$, and therefore $|(t+S)\cap S|\geq|S|(|S|-1)/d$. ∎
In the second auxiliary claim, we identify a simple forbidden induced subgraph of $G$. For a positive integer $k$, let $H_{k}$ be the graph that is formed by a clique of size $2k$, and an additional vertex connected to exactly half of the vertices of the clique. We show that the smallest eigenvalue of $H_{k}$ is $\Omega(\sqrt{k})$. We remark that $H_{k}$ and its relatives have been studied in connection to the smallest eigenvalue problem for a long time, see e.g. [54].
@@
Let $V(H_{k})=X\cup Y\cup\{x_{0}\}$, where $X\cup Y$ is a clique of size $2k$, and $X$ is the neighbourhood of $x_{0}$. Then we have $|X|=|Y|=k$. Let $B$ be the adjacency matrix of $H_{k}$ and let $v\in\mathbb{R}^{V(H_{k})}$ be the vector defined as
$v(x)=\begin{cases}\frac{1}{\sqrt{2}}&\mbox{if }x=x_{0}\\ -\frac{1}{2\sqrt{k}}&\mbox{if }x\in X\\ \frac{1}{2\sqrt{k}}&\mbox{if }y\in Y.\end{cases}$
Then $\|v\|_{2}=1$ and thus
$\mu\leq v^{T}Bv=2\sum_{xy\in E(G)}v(x)v(y)=-\sqrt{\frac{{k}}{2}}-\frac{1}{2}<-\sqrt{\frac{k}{2}}.\hfill\qed$
By Claim 3.3 and Cauchy’s interlacing theorem (cf. Section 4), $G$ does not contain $H_{k}$ as an induced subgraph for $k=2d^{2\gamma}=2d^{1/5-2\varepsilon}$. As $G[S]$ is a clique with $|S|>2k$, each vertex of $G$ sends either at most $k$ edges to $S$, or at least |S|k|S|-k edges. We prove that every vertex in $G$ must send at least |S|k|S|-k edges to $S$. This easily leads to a contradiction for $n$ sufficiently large: this implies that there are at least $(n-|S|)(|S|-k)\geq\frac{n}{2}\cdot\frac{|S|}{2}>d|S|$ edges with an endpoint in $S$, contradicting that $G$ is $d$-regular.
@@
Every $v\in V(G)$ sends at least |S|k|S|-k edges to $S$.
@@
Let $t\in\mathbb{Z}/n\mathbb{Z}$ be a non-zero element such that $|(t+S)\cap S|\geq\frac{|S|(|S|-1)}{d}$, whose existence is guaranteed by Claim 3.2. We prove by induction on $\ell$ that every vertex of $\ell t+S$ sends at least |S|k|S|-k edges to $S$. As every vertex $v\in V(G)$ is contained in some $\ell t+S$, this finishes the proof. The base case $\ell=0$ is trivial, so let $\ell\geq 1$. By our induction hypothesis and translation invariance, every vertex $v\in\ell t+S$ sends at least |S|k|S|-k edges to $t+S$. But then $v$ sends at least
$|S\cap(t+S)|-k=\frac{|S|^{2}}{2d}-k\geq\frac{1}{2}d^{1/5+8\varepsilon}-2d^{1/5-2\varepsilon}>2d^{1/5-2\varepsilon}=k$
edges to $S\cap(t+S)$, and in particular, more than $k$ edges to $S$. Therefore, as $G$ contains no induced copy of $H_{k}$, $v$ must send at least |S|k|S|-k edges to $S$, and we are done. ∎
@@ 4 Preliminaries
We recall some basic facts and standard notation from linear algebra and graph theory. The edge density of an $n$-vertex graph $G$ is $m/\binom{n}{2}$, where $m=e(G)$ is the number of edges. Given a subset $U$ of the vertices, $G[U]$ denotes the subgraph of $G$ induced on vertex set $U$. Also, if $V\subset V(G)$ is disjoint from $U$, then $G[U,V]$ is the bipartite subgraph of $V(G)$ induced between $U$ and $V$. The complement of $G$ is denoted by $\overline{G}$. The maximum degree of $G$ is denoted by $\Delta(G)$, and the average degree by $d(G)$. We will often identify the set of vertices of $G$ with $n]=\{1,2,\dots,n\}$.
@@
Given an $n\times n$ real symmetric matrix $M$, we denote by $\lambda_{1}(M)\geq\dots\geq\lambda_{n}(M)$ the eigenvalues of $M$ with multiplicity. If $G$ is an $n$-vertex graph whose adjacency matrix is $A$, then we denote by $\lambda_{i}=\lambda_{i}(A)$ the eigenvalues of $A$, sometimes also calling them the eigenvalues of $G$. We also denote by $v_{1},\dots,v_{n}$ a corresponding orthonormal basis of eigenvectors (all vectors in this paper will be column vectors by default). By the Perron–Frobenius theorem, we may take $v_{1}$ to be a vector with non-negative entries, which we call the principal eigenvector of $A$. Furthermore, the corresponding eigenvalue satisfies $d(G)\leq\lambda_{1}\leq\Delta(G)$. See the survey [28] as a general reference on the principal eigenvector.
An important and useful fact about spectra of graphs is Cauchy’s interlacing theorem. In the case of graphs, it states that if $G$ is an $n$-vertex graph with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$ and $G^{\prime}\subseteq G$ is an induced subgraph on n1n-1 vertices with eigenvalues $\mu_{1}\geq\dots\geq\mu_{n-1}$, then we have
$\lambda_{1}\geq\mu_{1}\geq\lambda_{2}\geq\mu_{2}\geq\dots\geq\mu_{n-1}\geq\lambda_{n}.$
For a proof of this result, see e.g. [42]. Crucially, this implies that if $G$ is a graph with smallest eigenvalue $\lambda_{n}$ and $G^{\prime}\subseteq G$ is an induced subgraph of $G$ with smallest eigenvalue $\mu_{k}$, then $\mu_{k}\geq\lambda_{n}$.
@@
Here, we use the Hadamard product for vectors: for $u,v\in\mathbb{R}^{n}$, their Hadamard product vector $u\circ v\in\mathbb{R}^{n}$ is defined by $(u\circ v)(i):=u(i)v(i)$ for all $i\in[n]$.
Throughout our proofs, we omit the use of floors and ceilings whenever they are not crucial.
@@ 5 Spectral lower bounds for the surplus
In this section, we present bounds on the MaxCut of a graph in terms of its spectrum. These inequalities are crucial for transferring our results for the smallest eigenvalue to the MaxCut setting.
@@
For an $n$-vertex graph $G$ with smallest eigenvalue $\lambda_{n}$, we have $\operatorname{surp}(G)\leq|\lambda_{n}|n/4$.
@@
Note that $x^{T}Ax\leq|\lambda_{n}|\|x\|_{2}^{2}$ for every vector $x\in\mathbb{R}^{n}$, so $\operatorname{surp}(G)\leq\frac{1}{4}|\lambda_{n}|\sqrt{n}^{2}=|\lambda_{n}|n/4$. ∎
The key ingredient of the above proof is the relation $\operatorname{surp}(G)=\frac{1}{4}\max_{x\in[-1,1]^{n}}-x^{T}Ax$. This can also be written as $\operatorname{surp}(G)=\frac{1}{4}\max_{x\in[-1,1]^{n}}\langle-A,xx^{T}\rangle$, where we observe that $xx^{T}$ is a positive-semidefinite matrix with diagonal entries bounded by $1$. Based on this, we define the semidefinite relaxation of the surplus as follows. Given an $n$-vertex graph $G$ with adjacency matrix $A$, define
@@
$\|v\|_{\infty}\leq\frac{\sqrt{n}}{|\lambda|}.$
@@
where we used that $\sum_{i=1}^{n}v_{1}(i)^{2}=1$. Hence, $|\lambda||v(b)|\leq\sqrt{n}$, i.e. $|v(b)|\leq\sqrt{n}/|\lambda|$. ∎
@@
Let $G$ be a graph on $n>10$ vertices, whose complement has edge density $p\leq 1/10$ and maximum degree $\bar{\Delta}=\Delta(\overline{G})$. If $v_{1}$ is the principal eigenvector of $G$, then for each $i\in[n]$ we have
$\frac{1-3\bar{\Delta}/n}{\sqrt{n}}\leq v_{1}(i)\leq\frac{1+2p+2/n}{\sqrt{n}}.$
@@
Let d=d(G)=(1p)(n1)d=d(G)=(1-p)(n-1) be the average degree of $G$, and recall that $\lambda_{1}\geq d$. By Lemma˜5.4,
$v_{1}(i)\leq\frac{\sqrt{n}}{\lambda_{1}}\leq\frac{\sqrt{n}}{d}=\frac{\sqrt{n}}{(1-p)(n-1)}\leq\frac{1+2p+2/n}{\sqrt{n}}.$
In the last inequality, we used that $p<1/10$ and $n>10$. To prove the lower bound, we may assume that $\bar{\Delta}\geq 1$. Observe that
$1=\sum_{i=1}^{n}v_{1}(k)^{2}\leq\|v_{1}\|_{\infty}\sum_{k=1}^{n}v_{1}(k)\leq\frac{\sqrt{n}}{|\lambda_{1}|}\sum_{k=1}^{n}v_{1}(k),$
which implies that $\sum_{k=1}^{n}v_{1}(k)\geq\frac{\lambda_{1}}{\sqrt{n}}$. Then, using the identity $Av_{1}=\lambda_{1}v_{1}$,
$\lambda_{1}v_{1}(i)=\sum_{k\sim i}v_{1}(k)\geq\sum_{k=1}^{n}v_{1}(k)-(\bar{\Delta}+1)\|v_{1}\|_{\infty}\geq\frac{\lambda_{1}}{\sqrt{n}}-\frac{(\bar{\Delta}+1)\sqrt{n}}{\lambda_{1}}=\frac{\lambda_{1}}{\sqrt{n}}\Big(1-\frac{(\bar{\Delta}+1)n}{\lambda_{1}^{2}}\Big)\!\!\geq\frac{\lambda_{1}}{\sqrt{n}}\Big(1-\frac{3\bar{\Delta}}{n}\Big),$
where we used $\lambda_{1}^{2}\geq d^{2}\geq(9/10)^{2}(n-1)^{2}\geq(2/3)n^{2}$ in the end. Cancelling $\lambda_{1}$ gives $v_{1}(i)\geq\frac{1-3\bar{\Delta}/n}{\sqrt{n}}$. ∎
Proof of Lemma˜5.3..
@@
(i) Let $X=\sum_{\lambda_{i}<0}v_{i}v_{i}^{T}$. Then $X$ is positive semidefinite, and as $v_{1},\dots,v_{n}$ is an orthonormal basis,
@@
(iii) Let $\beta=\frac{1}{120(\Delta^{*}+1)}$, and $X=\beta\sum_{\lambda_{i}<0}\lambda_{i}^{2}v_{i}v_{i}^{T}$. Then $X$ is positive semidefinite. It is enough to prove that the diagonal entries of $X$ are bounded by $1$, as then $\operatorname{surp}^{*}(G)\geq-\langle A,X\rangle=\beta\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}$.
First, consider the (easier) case $\Delta^{*}=\Delta(G)$. Observe that
@@
is positive semidefinite, so the diagonal entries of $X$ are dominated by those of $\beta A^{2}$, which are simply the degrees in $G$. So, $X_{i,i}\leq\beta(A^{2})_{i,i}\leq\beta\Delta\leq 1$, as claimed.
Next, consider the case $\Delta^{*}=\Delta(\overline{G})$. We may assume that the edge density of $\overline{G}$ is less than $1/10$, otherwise $\Delta^{*}=\Omega(n)$, and the previous case implies $\operatorname{surp}^{*}(G)=\Omega\big(\frac{1}{n}\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}\big)$. To show that $X_{i,i}\leq 1$, we analyse the matrix $B=A-\lambda_{1}v_{1}v_{1}^{T}$. Since
@@
we have that $\beta B^{2}-X$ is positive semidefinite. This means $X_{i,i}\leq(\beta B^{2})_{i,i}$ for every $i\in[n]$. Therefore, it is enough to show that $(B^{2})_{i,i}\leq 1/\beta=120(\Delta^{*}+1)$.
To this end, we first bound the entries of $B$. We denote by $p$ the density of $\overline{G}$, and observe that $p\leq\frac{\Delta^{*}n/2}{\binom{n}{2}}=\frac{\Delta^{*}}{n-1}$. Then, Lemma˜5.5 implies that for any $i,j\in[n]$ we have
$1-\frac{7(\Delta^{*}+1)}{n}\leq(n-1)(1-p)\left(\frac{1-3\Delta^{*}/n}{\sqrt{n}}\right)^{2}\leq\lambda_{1}v_{1}(i)v_{1}(j)\leq n\left(\frac{1+2p+2/n}{\sqrt{n}}\right)^{2}\leq 1+\frac{7(\Delta^{*}+1)}{n}.$
Therefore, for every $i,j\in[n]$, if $ij\in E(G)$ and $A_{i,j}=1$, then $|B_{i,j}|\leq\frac{7(\Delta^{*}+1)}{n}$. Otherwise, we have $|B_{i,j}|\leq 1+7(\Delta^{*}+1)/n\leq 8$. From this, we have
$(B^{2})_{i,i}=\sum_{j=1}^{n}(B_{i,j})^{2}\leq 64(\Delta^{*}+1)+n\frac{49(\Delta^{*}+1)^{2}}{n^{2}}\leq 120(\Delta^{*}+1).$
@@
Note that the first inequality is the combination of (i) and (iii), while the second one is simply the Cauchy–Schwartz inequality applied to the sequences $(|\lambda_{i}|^{3})_{\lambda_{i}<0}$ and $(|\lambda_{i}|)_{\lambda_{i}<0}$. Taking square roots then proves (ii). ∎
@@
Write $A^{\prime}$ for the adjacency matrix of $G^{\prime}$, and let $X^{\prime}\in\mathbb{R}^{V(G^{\prime})\times V(G^{\prime})}$ be a matrix such that $X^{\prime}$ is positive semidefinite, $X^{\prime}_{i,i}\leq 1$ for every $i\in V(G^{\prime})$, and $\operatorname{surp}^{*}(G^{\prime})=-\langle A^{\prime},X^{\prime}\rangle$. Then, write $A$ for the adjacency matrix of $G$ and let $X\in\mathbb{R}^{V(G)\times V(G)}$ be the matrix that agrees with $X^{\prime}$ on every entry $(x,y)\in V(G^{\prime})\times V(G^{\prime})$, and zero everywhere else. Then
$\operatorname{surp}^{*}(G)\geq-\langle A,X\rangle=-\langle A^{\prime},X^{\prime}\rangle=\operatorname{surp}^{*}(G^{\prime}).\hfill\qed$